1. Suppose you are hired by the Federal Reserve Bank of St. Louis as a financial analyst. Suppose also the tax rates and brackets are also given:

Federal Income Tax Brackets

Taxable Income

    Tax Rate

$0–9,700

10 percent

$9,701–$39,475

12 percent

$39,476–$84,200

22 percent

$84,201–$160,725

24 percent

$160,726–$204,100

$204,101–$510,300

32 percent

35 percent

Over $510,300

37 percent

  1. Calculate the total tax bill and average tax rate if your salary is $78,000
  2. Calculate the total tax bill and average tax rate if your salary is $180,000
Answer #1

a)Salary = 78000

Tax bill = (9700*10%)+(39475-9700*12%)+(78000-39475*22%)

= 970+3573+8475.5 = 13018.5

Average tax rate = 13018.5/78000*100 = 16.69%

b) Salary = 180000

Tax bill = (970)+(3573)+(84200-39475*22%)+(160725-84200*24%)+(180000-160725*32%)

= 970+3573+9839.5+18366+6168 = 38916.5

Average tax rate = 38916.5/180000*100 = 21.62%